OpenStudy (anonymous):

How could you eliminate the fractions to make solving the equation five twenty-thirds m plus twelve twenty-thirds equals twenty-one twenty-thirds easier? Divide both sides of the equation by five twenty-thirds. Divide both sides of the equation by twenty-three. Multiply both sides of the equation by twenty-three. Subtract twelve twenty-thirds from both sides of the equation.

OpenStudy (texaschic101):

5/23m + 12/23 = 21/23 --- multiply the equation by 23 5m + 12 = 21 5m = 21 - 12 5m = 9 m = 9/5

OpenStudy (texaschic101):

if you multiply the equation by the LCD, then it will get rid of the fractions